OpenStudy (anonymous):

The diagonals of a rectangle have lengths 4 + 2x and 6x – 20. x =

OpenStudy (anonymous):

as diagonals of a rectangle are equal 4 + 2x = 6x - 20 24 = 4x 6 = x

OpenStudy (anonymous):

since diagonals of a rectangle r equal 4+2x = 6x - 20 4 + 20 = 6x - 2x 24/4 = x 6 = x

OpenStudy (anonymous):

The diagonals of a rectangle are equal, so 4 + 2x = 6x - 20. So x = 6.

OpenStudy (anonymous):

AWSOME!